# An innovative clay metaBrick-based motif to enhance thermal and acoustic insulation

### v1

@@ 3.1 Experimental and numerical setups
The experimental setups used in this study have been mounted to investigate the acoustic and thermal behaviors of the novel metaBrick and compared to a standard brick. The schematics of these setups are illustrated in Fig. 7(a) and (b). Thus, these setups were composed of a double steel box, as depicted in Fig. 7(c) and (d), in which one side is left open to be filled with the brick-based walls. The walls were constructed using the metaBrick and the standard brick.
@@
The sound pressure level in the box is measured with a calibrated measurement microphone. The microphone was positioned in a fixed location throughout the experiment and numerical test because the sound pressure level in the box varies from one point to another, even in an ideal diffuse sound field. For that, we have conserved the same point for both experimental and numerical tests. The collected data, which represents the pressure in function of time, is processed with the fast Fourier transformation to the amplitude of the sound with the expression $amplitude=20*\log_{10}(p)$. The results of this study are illustrated in Fig. 8.
In the thermal study, a heat source was used as the excitation source. Both walls were exposed to an outer temperature of $96$$\mathrm{\SIUnitSymbolDegree}\mathrm{C}$ for $600$ minutes in an ambient temperature of $25$$\mathrm{\SIUnitSymbolDegree}\mathrm{C}$, and the temperature fluctuations were measured using infrared thermometers that were positioned at a fixed distance from the wall surface on both sides, the inner side and the outer side. The thermometers were used to measure the temperature evolution of the wall at different time intervals. The results of this study are illustrated in Fig. 9.
@@ 3.2 Results
The results obtained from the acoustic analysis provided valuable insights into the effectiveness of the metaBrick design in attenuating sound, as shown in Fig. 8. The analysis of the four signals obtained, namely the excitation chirp signal (in violet), the measured excitation signal without walls (in blue), steel box sound transmission (in pink), standard brick sound transmission (in black), and metaBrick signal transmission (in red), enabled a comparison of their respective acoustic behaviors. The numerical and experimental results are presented in Fig. 8(a) and (b), respectively. It is clearly seen that the numerical results and the experimental results are in good agreement; in what follows, we interpret both results together.
@@
In contrast, the metaBrick transmission signal showed a very weak transmission curve in contrast to the measured excitation signal and the transmission of the standard brick. The amelioration is based on the presence of the slits in the metaBrick, which had excellent performance in the attenuation of the sound at low frequencies. The transmission loss for the wall built on the metaBrick up to $40$ $\mathrm{d}\mathrm{B}$ compared to the measured excitation signal and $20$ $\mathrm{d}\mathrm{B}$ compared to the transmission loss generated by the standard brick.
@@
To elaborate further, the thermal study involved measuring the temperature variation inside and outside the iron box over a period of 600 minutes while exposed to a temperature of $90$$\mathrm{\SIUnitSymbolDegree}\mathrm{C}$. Inner and outer temperature evolution is measured for both numerical and experimental studies. The results of this study are presented in Fig. 9. It is clearly seen that the numerical and the experimental results are in good agreement.
@@ 4 Discussion
We conducted both numerical and experimental compression tests on the metaBrick and standard clay brick. Fig. 10 shows the numerical compression stress-strain curve and the mechanical resistance of bricks during compression, which is measured using a compression testing machine. Besides, experimentally we examined four samples of each type of brick and determined their compressive strength in $\mathrm{M}\mathrm{P}\mathrm{a}$. The results of the study are summarized in Table 2.
Figure 10: The numerical compression stress-strain curve is in line, and the experimental compressive strength is in dashed points for both standard brick and metabrick.
The standard brick has an average compressive strength of around $19$ $\mathrm{M}\mathrm{P}\mathrm{a}$, whereas the metaBrick has an average compressive strength of about $12$ $\mathrm{M}\mathrm{P}\mathrm{a}$. Although metaBrick has lost strength, it still possesses a relatively high compressive strength that is more than enough for the building sector ($7.5$ $\mathrm{M}\mathrm{P}\mathrm{a}$) gencel2015characteristics . The reason for the strength loss is due to the existence of slits in metaBrick, which causes the compressive strength to decrease by approximately $7$ $\mathrm{M}\mathrm{P}\mathrm{a}$, which is equivalent to $33\%$ when compared with standard brick.
Table 2: Compressive strength of bricks.
Type of brick Sample Compressive strength ($\mathrm{M}\mathrm{P}\mathrm{a}$) Average compressive strength
Standard clay brick 1 19.5 19.37
@@ 5 Conclusions
In conclusion, the present study has investigated the acoustic and thermal behaviors of a novel metamaterial design named metaBrick. Numerical simulations and experimental measurements have demonstrated that the presence of slits in alveoli in the metaBrick structure creates Helmholtz resonators, which can effectively attenuate both sound and heat. The findings of this study indicate that metaBrick has a strong potential for improving insulation performance in building materials with a compressive strength that is high enough for the construction sector. The results also indicate that further research is warranted to explore the use of metamaterials in multi-functional applications to enhance building performance and energy efficiency. Overall, this study has demonstrated the importance of developing new and innovative building materials to meet the increasing demand for sustainable and energy-efficient buildings. Metamaterials, such as the metaBrick design, have shown potential for making significant contributions to this effort.
Acknowledgements
We would like to thank the Moroccan Ministry of Higher Education, Scientific Research and Innovation and the OCP Foundation who funded this work through the APRD research program.

### v2

@@ 3.1 Experimental and numerical setups
The experimental setups used in this study have been mounted to investigate the acoustic and thermal behaviors of the novel metaBrick and compared to a standard brick. The schematics of these setups are illustrated in Figure 6(a) and (b). Thus, these setups were composed of a double steel box, as depicted in Figure 6(c) and (d), in which one side is left open to be filled with the brick-based walls. The walls were constructed using the metaBrick and the standard brick.
@@
The sound pressure level in the box is measured with a calibrated measurement microphone. The microphone was positioned in a fixed location throughout the experiment and numerical test because the sound pressure level in the box varies from one point to another, even in an ideal diffuse sound field. For that, we have conserved the same point for both experimental and numerical tests. The collected data, which represents the pressure in function of time, is processed with the fast Fourier transformation to the amplitude of the sound with the expression $amplitude=20*\log_{10}(p)$. The results of this study are illustrated in Figure 7.
In the thermal study, a heat source was used as the excitation source. Both walls were exposed to an outer temperature of $96$$\mathrm{\SIUnitSymbolDegree}\mathrm{C}$ for $600$ minutes in an ambient temperature of $25$$\mathrm{\SIUnitSymbolDegree}\mathrm{C}$, and the temperature fluctuations were measured using an calibrated sensors (infrared thermometers) that were positioned at a fixed distance from the wall surface on both sides, the inner side and the outer side. The thermometers were used to measure the temperature evolution of the wall at different time intervals. The results of this study are illustrated in Figure 8.
@@ 3.2 Results
The results obtained from the acoustic analysis provided valuable insights into the effectiveness of the metaBrick design in attenuating sound, as shown in Figure 7. The analysis of the four signals obtained, namely the excitation chirp signal (in violet), the measured excitation signal without walls (in blue), steel box sound transmission (in pink), standard brick sound transmission (in black), and metaBrick signal transmission (in red), enabled a comparison of their respective acoustic behaviors. The numerical and experimental results are presented in Figure 7(a) and (b), respectively. It is clearly seen that the numerical results and the experimental results are in good agreement; in what follows, we interpret both results together.
@@
In contrast, the transmission signal of the metabrick exhibited a much weaker transmission curve compared to both the measured excitation signal and the transmission of the standard brick. This improvement is primarily attributed to the presence of the slits in the metabrick, which performed exceptionally well in attenuating sound at low frequencies. The transmission loss for the wall constructed with the metabrick reached up to $40$ $\mathrm{d}\mathrm{B}$ compared to the measured excitation signal and $20$ $\mathrm{d}\mathrm{B}$ compared to the transmission loss of the standard brick. However, both the numerical and experimental results show some discrepancies, which may be due to differences in the material properties used in the simulations and the real-world scenario. For the numerical models, it is important to note that we assumed all materials to be linear isotropic elastic, which does not reflect real-world conditions. In practice, additional factors such as porosity, viscosity, and other material characteristics can influence the results.
@@
To elaborate further, the thermal study involved measuring the temperature variation inside and outside the iron box over a period of 600 minutes while exposed to a temperature of $90$$\mathrm{\SIUnitSymbolDegree}\mathrm{C}$. Inner and outer temperature evolution is measured for both numerical and experimental studies. The results of this study are presented in Figure 8. It is clearly seen that the numerical and the experimental results are in good agreement.
@@ 4 Discussion
We conducted experimental compression tests on both the metabrick and the standard clay brick. Four samples of each brick type were tested to determine their compressive strength, expressed in $\mathrm{M}\mathrm{P}\mathrm{a}$. The results are summarized in Table 2.
Table 2: Experimental compressive strength of bricks.
Type of brick Sample Compressive strength ($\mathrm{M}\mathrm{P}\mathrm{a}$)
1 19.5
@@
5 12
The standard brick has an average compressive strength of around $19$ $\mathrm{M}\mathrm{P}\mathrm{a}$, whereas the metaBrick has an average compressive strength of about $12$ $\mathrm{M}\mathrm{P}\mathrm{a}$. Although metaBrick has lost strength, it still possesses a relatively high compressive strength that is more than enough for the building sector ($7.5$ $\mathrm{M}\mathrm{P}\mathrm{a}$) [55]. The reason for the strength loss is due to the existence of slits in metaBrick, which causes the compressive strength to decrease by approximately $7$ $\mathrm{M}\mathrm{P}\mathrm{a}$, which is equivalent to $33\%$ when compared with standard brick.
@@ 5 Conclusions
In conclusion, the present study has investigated the acoustic and thermal behaviors of a metamaterial design named metaBrick. Numerical simulations and experimental measurements have demonstrated that the presence of slits in alveoli in the metaBrick structure creates Helmholtz resonators, which can effectively attenuate both sound and heat. The findings of this study indicate that metaBrick has a strong potential for improving insulation performance in building materials with a compressive strength that is high enough for the construction sector. The results also indicate that further research is warranted to explore the use of metamaterials in multi-functional applications to enhance building performance and energy efficiency. Overall, this study has demonstrated the importance of developing new and innovative building materials to meet the increasing demand for sustainable and energy-efficient buildings. Metamaterials, such as the metaBrick design, have shown potential for making significant contributions to this effort.
Acknowledgments
We would like to thank the Moroccan Ministry of Higher Education, Scientific Research and Innovation and the OCP Foundation who funded this work through the APRD research program. We acknowledge the support of the ANR PNanoBot project (contract ”ANR-21-CE33-0015”).
Appendix A: Sound transmission as a function of slit width variation.
Figure 9: Sound transmission as a function of slit width variation.
Appendix B: Displacement distribution in the metabrick.
Figure 10: Screenshots of the displacement modulus of the metaBrick. (a) At a frequency of 500 Hz. (b) At a frequency of 671 Hz. (c) At a frequency of 793 Hz. (d) At a frequency of 1500 Hz.